OpenStudy (anonymous):

write the expression with positive exponents only then simplify if possible... x2y-3 _______ z-4

OpenStudy (anonymous):

\[\frac{x^2y^{-3}}{z^{-4}}=\frac{x^2z^4}{y^3}\]

OpenStudy (anonymous):

with a negative exponent if it is up bring it down, and if it is down bring it up.

OpenStudy (anonymous):

THANK YOU !!